# Short-Context dominance: How Much Local Context Natural Language Actually Needs?

### v1

@@ 1 Introduction
When prompted to continue a piece of text, how much preceding context do humans rely on? Do they focus on recent words and local coherence, or plan with a broader, narrative-wide perspective? For example, when writing a story, do they recall events from earlier chapters or rely mostly on recent developments? While difficult to study these questions rigorously in humans, large language models (LLMs) offer a tractable experimental analogue. This leads us to ask: How far back must a model look to predict the next token accurately? Although modern transformer-based LLMs can attend to thousands of tokens (longformer; transformerXL), it remains unclear and unquantified how often that capacity is used at inference time, and how often predictions depend on distant information or primarily on local spans.
@@
Figure 2: Distribution of MCL: Minimum context window needed to confidently predict the next token across sampled sequences, six datasets, and three LLMs. $\hat{b}$ denotes the slope of the log-log fit. Blue/green distinguishes between long ($\geq$ 6k tokens) and short ($\geq$ 1k tokens) documents, respectively.
@@
Post-hoc short-context dominance correction. With a long-context detector at hand, we finally test a hypothesis that short-context dominance induces a corresponding bias in LLMs themselves: Since sequences predominantly require only short context, models are implicitly trained on distributions heavily skewed toward local dependencies, biasing them toward common completions and filler words predictable from short local context, hurting prediction for long-context sequences (sharma2023truth; CoherenceBoosting; CADBoost). We validate such a short-context bias by applying TaBoo (Targeted Boosting ), an intuitive decoding algorithm that counteracts the skew: for sequences that we detect as being long-context, TaBoo identifies and boosts tokens that are long-context-relevant. On Q&A datasets with inherent long-context dependencies, TaBoo consistently outperforms vanilla nucleus sampling and competitive logit-adjustment methods across model architectures.
@@ 2 Related Work
Long context utilization analysis. Prior studies demonstrate that LMs often underutilize available context (khandelwal2018sharpfuzzy; sun2021longcontext; liu2023lostmiddlelanguagemodels). These findings have inspired algorithmic interventions at both system and data levels (foundinthemiddle; borgeaud2022improving; izacard2022atlasfewshot; checnLDAMlongcontext; selfcite). A more closely related recent study by PerplexityLongCtx provides methodology for detecting tokens with long context relevance and emphasizing them upon evaluation and fine-tuning. Our work differs by providing systematic quantification of minimal context requirements from the perspective of natural language properties themselves, introducing practical detection methods that operate without ground-truth tokens (a key distinction to PerplexityLongCtx), and demonstrating targeted inference-time interventions.
Contrastive decoding. Several inference-time methods address hallucination in long-context generation through contrastive approaches (li2023contrastivedecoding; zhao2024contextual; liu2021dexperts). Building on (diversitypromoting2016; brown2020language), CoherenceBoosting; CADBoost reweight distributions by contrasting long vs. short contexts. CAD (CADBoost) is most closely related to our TaBoo algorithm, but differs fundamentally in both motivation and implementation. While CAD uniformly adjusts output probabilities for all sequences and tokens through contrastive considerations, our algorithm stems from the short-context dominance hypothesis and applies targeted adjustments to long-context sequences and their long-range-relevant tokens. vanderpoel2022mutual apply entropy-based selection for contrastive decoding, but their method differs again in both theoretical motivation and technical implementation. Beyond distinctions specific to TaBoo, we systematically study minimal context length requirements in natural language and our long-context detection methods could apply beyond inference-time correction.
n-gram models. Recent work has revisited n-gram models as complementary to neural language models (ngramisback; liu2025infinigramscalingunboundedngram; nguyenngramTransfomer). The success of such inherently short-range models (even liu2025infinigramscalingunboundedngram’s Infini-gram typically captures at most 32-token dependencies) provides indirect evidence toward short-context dominance in natural language. Our systematic quantification of minimal context requirements offers a principled explanation of this, supporting the broader thesis that majority of language understanding tasks require primarily local information.
@@ 3.2 Experimental Setup
Datasets. We experiment on the following natural-language documents. Short documents: Reddit Writing Prompts (fan-etal-2018-hierarchical), CNN/DailyMail News Articles (hermann2015teaching; nallapati2016abstractive) and Wikipedia articles from WikiText-103 (merity2016pointer). Long documents: U.S. Government Reports from GovReport (huang2021efficient), Meeting Transcripts from QMSum qmsum and story Book Chapters BookSum (booksum). These appear in the LongBench and LongEval (bai2024longbenchbilingualmultitaskbenchmark; longeval) benchmarks and are believed to have inherent long-context qualities. We further experiment with documents of different languages using (wikimedia) and datasets specialized for math and medical domains (Appx.  C.2).
LLM Oracles. We evaluate three open-weight models: LLaMA-3-8B (grattafiori2024llama3herdmodels), Mistral-7B-Instruct (v0.2) (jiang2023mistral7b) and Qwen2-7B (yang2024qwen2technicalreport). We assume that these models are sufficiently capable to exhibit reliable performance on next-token prediction and question answering tasks. Recall also that we isolate sequences that the models confidently predict the next token. All experiments are performed on a V100 Nvidia GPU with 32GB of memory.
@@
MCL Algorithm. To determine the MCL as per Defn. 1, we evaluate a model’s predictions using increasing prefix sizes $l\in\{32,48,64,\dots,|\mathbf{s}|\}$, starting from 32 tokens and incrementing by 16. For the longer documents, we start from 32 and increment by 64. Starting from 32 tokens is motivated by prior work suggesting this length captures local context beyond classical n-gram statistics (CoherenceBoosting; liu2025infinigramscalingunboundedngram; PerplexityLongCtx). For each window size, we evaluate the model’s output distribution and stop once it confidently predicts the ground-truth next token. In practice, we provide the full input to preserve positional encoding and simulate truncated contexts via attention masking.
@@ 3.3 Results and Discussion
Validating the short-context dominance hypothesis carries implications for both pretraining and evaluation: Since most predictions require only highly localized information, then standard training objectives and perplexity-based evaluations are inherently skewed toward short-range dependencies, potentially obscuring true progress on long-context reasoning. This helps explain prior findings that only small fraction of tokens benefit from contexts larger than 2K (sun2021longcontext). It also reinforces recent critiques (PerplexityLongCtx) that token-level perplexity is an insufficient metric for long-context evaluation—even on datasets with genuine long context dependencies, such as Government Reports.
@@ 4 Distributional Awareness
MCL evaluates whether the ground-truth next token from the dataset can be predicted with a shorter prefix. Intuitively, natural language often permits multiple valid next tokens, and models may assign high probability to plausible alternatives that differ from the ground-truth token in the dataset. Definition 1 is also limited to greedy decoding, while popular natural language generation methods often rely on sampling strategies that draw from multiple probable tokens (holtzman2020curious; basu2021mirostatneuraltextdecoding; zhu2024adaptive). These considerations (see also Appx. D.1) motivate us to introduce here a more flexible MCL formulation.
@@ 4.1 Distribution Aware MCL
Consider decoding strategy $\phi$ that modifies the model’s raw probability distribution. For example, nucleus sampling (holtzman2020curious) selects a subset of tokens whose cumulative probability mass reaches threshold $p$ (e.g. $p=0.9$), then produces a renormalized distribution $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$ with support on this subset and zero probability to tokens otherwise. More generally, our framework accommodates any decoding strategy $\phi$ that produces a probability distribution $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$. We introduce DaMCL to find the smallest prefix $\mathbf{s}_{[-\ell:]}$ such that $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s}_{[-\ell:]})$ is sufficiently similar to $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$.
@@
While this definition accommodates any distance metric $\mathcal{M}$, we specifically use the Jensen-Shannon Distance (JSD) throughout our experiments. For distributions $\mathbf{p}_{1},\mathbf{p}_{2}$ in the $|\mathcal{V}|$-dimensional simplex, JSD is defined as $\operatorname{JSD}(\mathbf{p}_{1}\,;\,\mathbf{p}_{2}):=\sqrt{\frac{1}{2}\mathrm{KL}(\mathbf{p}_{1}\|\mathbf{q})+\frac{1}{2}\mathrm{KL}(\mathbf{p}_{2}\|\mathbf{q})}$, where $\mathbf{q}:=\frac{1}{2}(\mathbf{p}_{1}+\mathbf{p}_{2})$, and $\mathrm{KL}(\mathbf{p}_{1}\|\mathbf{p}_{2}):=\sum_{t\in\mathcal{V}}[\mathbf{p}_{1}]_{t}\log\frac{[\mathbf{p}_{1}]_{t}}{[\mathbf{p}_{2}]_{t}}$ is the Kullback-Leibler divergence. We choose JSD because it is a proper distance metric satisfying the triangle inequality and is widely employed in knowledge distillation (miniLLM_KD) and language model evaluation (tailoringLLMGen). The above definition of DaMCL (1) relies on measuring the difference between distributions as opposed to a single token, and, (2) does not require ground-truth information regarding the actual next token of sequence $\mathbf{s}$.
@@ 4.2 Experimental Setup
DaMCL evaluation. We evaluate DaMCL over several decoding strategies: Top-$K$ sampling (K=1K{=}1 for greedy) with $K\in[10]$ (radford2019language; fan-etal-2018-hierarchical), nucleus sampling with p=0.9p{=}0.9 (holtzman2020curious), and adaptive sampling with $\epsilon{=}0.001$ (zhu2024adaptive). We evaluate two threshold values $\operatorname{JSD}\leq 0.1$ and $\operatorname{JSD}\leq 0.2$ to study the effect of varying similarity criteria.
@@ 4.3 Results and Discussion
When enforcing stricter similarity requirements ($\operatorname{JSD}\leq 0.1$) we find the distribution becomes flatter and even experiences a U-shaped structure. This suggests that a larger subset of tokens either resolve with short contexts or require nearly the full input to reach distributional convergence. We further analyze this in Appx. D.3 and discover that this bimodality is mainly attributed to the shorter sequences. Indeed, we find that the long documents, for which all sequences have more than 6k tokens, do not experience the bimodality to the same degree.
@@ 5.1 Distributional-aware long-context sequence detection
For concreteness, unless otherwise stated, we fix $\phi$ to nucleus sampling with parameter $p=0.9$ (holtzman2020curious). (Ablation results on different metrics and $p$ values in Appx E.2.) Finally, we fix the prefix length to $32$ based on our findings in Sec. 3 and consistent with liu2025infinigramscalingunboundedngram.
Controlled validation. To demonstrate that LSDS can effectively detect long-context sequences, we conduct a controlled needle-in-a-haystack experiment adapted from KamradtNIAH2023. We create short and long queries, such that only in the former the answer appears in the final 32 tokens; see Appx. E.1.1 for details and example illustration. Figure 5 shows the resulting LSDS distributions for Mistral-7B-Instruct (v0.2). We observe a clear separation between categories despite no ground-truth next-token information: short-context sequences concentrate at low LSDS values ($\leq 0.4$), while long-context sequences dominate high LSDS values ($\geq 0.7$). Similar patterns hold for LLaMA-3-8B/Qwen2-7B (see Appx. E.1.1).
@@ 5.2 Evaluation on Natural Text
In natural text, determining the true context dependency requires an oracle with access to the actual next token. We employ two such oracles to establish ground-truth short-context vs long-context labels for sequences. (1) MCL Oracle: For $\mathbf{s}$ with next-token $t$, we classify $\mathbf{s}$ as long-context iff $\mathsf{MCL}\left(\mathbf{s}|t\right)\geq 32$ (see Defn. 1). (2) LSD Oracle: Following PerplexityLongCtx, we classify $\mathbf{s}$ as long-context iff $\operatorname{LSD}(\mathbf{s}|t)>2$ & $\operatorname{LCL}(\mathbf{s}|t)\geq-1$ (details in Appx. E.1.4 ).Both oracles require knowledge of ground-truth next token and thus cannot be deployed at inference time. We show that LSDS-based classification agrees well with these oracles despite not knowing the ground-truth.
Figure 6: Distribution of $\mathsf{LSDS}\left(\mathbf{s}\right)$ for short- and long-contexts on 2 GovReport (long doc) and Reddit Writing Prompts (short doc), pooled across three models. Illustrates a clear distinction between long and short contexts.
@@ 6 Long-Context Token Boosting
If so, we could favor generating such tokens when we detect a sequence requires long-context reasoning (which, as shown in the previous section, we can reliably identify). Here, we show this is possible and evaluate our method on text-based question answering, a standard benchmark for assessing long-context reasoning capabilities (liu2023lostmiddlelanguagemodels; longeval; longformer; bai2024longbenchbilingualmultitaskbenchmark).

### v2

@@ 1 Introduction
When prompted to continue a piece of text, how much preceding context do humans rely on? Do they focus on recent words and local coherence, or plan with a broader, narrative-wide perspective? For example, when writing a story, do they recall events from earlier chapters or rely mostly on recent developments? While difficult to study these questions rigorously in humans, large language models (LLMs) offer a tractable experimental analogue. This leads us to ask: How far back must a model look to predict the next token accurately? Although modern transformer-based LLMs can attend to thousands of tokens (Beltagy et al., 2020; Dai et al., 2019), it remains unclear and unquantified how often that capacity is used at inference time, and how often predictions depend on distant information or primarily on local spans.
@@
Figure 2: Distribution of MCL: Minimum context window needed to confidently predict the next token across sampled sequences, six datasets, and three LLMs. $\hat{b}$ denotes the slope of the log-log fit. Blue/green distinguishes between long ($\geq$ 7k tokens) and short ($\geq$ 1k tokens) documents, respectively.
@@
Post-hoc short-context dominance correction. With a long-context detector at hand, we finally test a hypothesis that short-context dominance induces a corresponding bias in LLMs themselves: Since sequences predominantly require only short context, models are implicitly trained on distributions heavily skewed toward local dependencies, biasing them toward common completions and filler words predictable from short local context, hurting prediction for long-context sequences (Sharma et al., 2023; Malkin et al., 2022; Duh et al., 2024). We validate such a short-context bias by applying TaBoo (Targeted Boosting ), an intuitive decoding algorithm that counteracts the skew: for sequences that we detect as being long-context, TaBoo identifies and boosts tokens that are long-context-relevant. On Q&A datasets with inherent long-context dependencies, TaBoo consistently outperforms vanilla nucleus sampling and competitive logit-adjustment methods across model architectures.
@@ 2 Related Work
Long context utilization analysis. Prior studies demonstrate that LMs often underutilize available context (Khandelwal et al., 2018; Sun et al., 2021; Liu et al., 2023). These findings have inspired algorithmic interventions at both system and data levels (Zhang et al., 2024; Borgeaud et al., 2022; Izacard et al., 2022; Chen et al., 2025; Chuang et al., 2025). In more closely related recent studies Fang et al. (2025) and Helm et al. (2025) provide methodology for detecting tokens with long context relevance and emphasizing them during fine-tuning and pretraining respectively, to improve long context performance. Our work differs by providing systematic quantification of minimal context requirements from the perspective of natural language properties themselves, introducing practical detection methods that operate without ground-truth tokens (a key distinction to Fang et al. (2025) and Helm et al. (2025)), and demonstrating targeted inference-time interventions.
Contrastive decoding. Several inference-time methods address hallucination in long-context generation through contrastive approaches (Li et al., 2023; Zhao et al., 2024; Liu et al., 2021). Building on (Li et al., 2016; Brown et al., 2020), Malkin et al. (2022); Duh et al. (2024) reweight distributions by contrasting long vs. short contexts. CAD (Duh et al., 2024) is most closely related to our TaBoo algorithm, but differs fundamentally in both motivation and implementation. While CAD uniformly adjusts output probabilities for all sequences and tokens through contrastive considerations, our algorithm stems from the short-context dominance hypothesis and applies targeted adjustments to long-context sequences and their long-range-relevant tokens. van der Poel et al. (2022) apply entropy-based selection for contrastive decoding, but their method differs again in both theoretical motivation and technical implementation. Beyond distinctions specific to TaBoo, we systematically study minimal context length requirements in natural language and our long-context detection methods could apply beyond inference-time correction.
n-gram models. Recent work has revisited n-gram models as complementary to neural language models (Li et al., 2022; Liu et al., 2025; Nguyen, 2024). The success of such inherently short-range models (even Liu et al. (2025)’s Infini-gram typically captures at most 32-token dependencies) provides indirect evidence toward short-context dominance in natural language. Our systematic quantification of minimal context requirements offers a principled explanation of this, supporting the broader thesis that majority of language understanding tasks require primarily local information.
@@ 3.2 Experimental Setup
Datasets. We experiment on the following natural-language documents. Short documents: Reddit Writing Prompts (Fan et al., 2018), CNN/DailyMail News Articles (Hermann et al., 2015; Nallapati et al., 2016) and Wikipedia articles from WikiText-103 (Merity et al., 2016). Long documents: U.S. Government Reports from GovReport (Huang et al., 2021), Meeting Transcripts from QMSum Zhong et al. (2021) and story Book Chapters BookSum (Kryściński et al., 2022). These appear in the LongBench and LongEval (Bai et al., 2024; Krishna et al., 2023) benchmarks and are believed to have inherent long-context qualities. We further experiment with documents of different languages using (Schwenk et al., 2019) and datasets specialized for math and medical domains (Appx.  C.2).
LLM Oracles. We evaluate three open-weight models: LLaMA-3-8B (Grattafiori et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) and Qwen2-7B (Yang et al., 2024). We assume that these models are sufficiently capable to exhibit reliable performance on next-token prediction and question answering tasks. Recall also that we isolate sequences that the models confidently predict the next token. All experiments are performed on a V100 Nvidia GPU with 32GB of memory.
@@
MCL Algorithm. To determine the MCL as per Defn. 1, we evaluate a model’s predictions using increasing prefix sizes $l\in\{32,48,64,\dots,|\mathbf{s}|\}$, starting from 32 tokens and incrementing by 16. For the longer documents, we start from 32 and increment by 64. Starting from 32 tokens is motivated by prior work suggesting this length captures local context beyond classical n-gram statistics (Malkin et al., 2022; Liu et al., 2025; Fang et al., 2025). For each window size, we evaluate the model’s output distribution and stop once it confidently predicts the ground-truth next token. In practice, we provide the full input to preserve positional encoding and simulate truncated contexts via attention masking.
@@ 3.3 Results and Discussion
Validating the short-context dominance hypothesis carries implications for both pretraining and evaluation: Since most predictions require only highly localized information, then standard training objectives and perplexity-based evaluations are inherently skewed toward short-range dependencies, potentially obscuring true progress on long-context reasoning. This helps explain prior findings that only small fraction of tokens benefit from contexts larger than 2K (Sun et al., 2021). It also reinforces recent critiques (Fang et al., 2025) that token-level perplexity is an insufficient metric for long-context evaluation—even on datasets with genuine long context dependencies, such as Government Reports. We provide additional results regarding the implications of a token’s position in word (Appx. C.3) and its part-of-speech label (Appx. C.4) has on its MCL distribution.
@@ 4 Distributional Awareness
MCL evaluates whether the ground-truth next token from the dataset can be predicted with a shorter prefix. Intuitively, natural language often permits multiple valid next tokens, and models may assign high probability to plausible alternatives that differ from the ground-truth token in the dataset. Definition 1 is also limited to greedy decoding, while popular natural language generation methods often rely on sampling strategies that draw from multiple probable tokens (Holtzman et al., 2020; Basu et al., 2021; Zhu et al., 2024). These considerations (see also Appx. 17) motivate us to introduce here a more flexible MCL formulation.
@@ 4.1 Distribution Aware MCL
Consider decoding strategy $\phi$ that modifies the model’s raw probability distribution. For example, nucleus sampling (Holtzman et al., 2020) selects a subset of tokens whose cumulative probability mass reaches threshold $p$ (e.g. $p=0.9$), then produces a renormalized distribution $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$ with support on this subset and zero probability to tokens otherwise. More generally, our framework accommodates any decoding strategy $\phi$ that produces a probability distribution $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$. We introduce DaMCL to find the smallest prefix $\mathbf{s}_{[-\ell:]}$ such that $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s}_{[-\ell:]})$ is sufficiently similar to $\operatorname{\mathbf{p}}_{\phi}(\mathbf{s})$.
@@
While this definition accommodates any distance metric $\mathcal{M}$, we specifically use the Jensen-Shannon Distance (JSD) throughout our experiments. For distributions $\mathbf{p}_{1},\mathbf{p}_{2}$ in the $|\mathcal{V}|$-dimensional simplex, JSD is defined as $\operatorname{JSD}(\mathbf{p}_{1}\,;\,\mathbf{p}_{2}):=\sqrt{\frac{1}{2}\mathrm{KL}(\mathbf{p}_{1}\|\mathbf{q})+\frac{1}{2}\mathrm{KL}(\mathbf{p}_{2}\|\mathbf{q})}$, where $\mathbf{q}:=\frac{1}{2}(\mathbf{p}_{1}+\mathbf{p}_{2})$, and $\mathrm{KL}(\mathbf{p}_{1}\|\mathbf{p}_{2}):=\sum_{t\in\mathcal{V}}[\mathbf{p}_{1}]_{t}\log\frac{[\mathbf{p}_{1}]_{t}}{[\mathbf{p}_{2}]_{t}}$ is the Kullback-Leibler divergence. We choose JSD because it is a proper distance metric satisfying the triangle inequality and is widely employed in knowledge distillation (Gu et al., 2024) and language model evaluation (Ji et al., 2023a). The above definition of DaMCL (1) relies on measuring the difference between distributions as opposed to a single token, and, (2) does not require ground-truth information regarding the actual next token of sequence $\mathbf{s}$.
@@ 4.2 Experimental Setup
DaMCL evaluation. We evaluate DaMCL over several decoding strategies: Top-$K$ sampling (K=1K{=}1 for greedy) with $K\in[10]$ (Radford et al., 2019; Fan et al., 2018), nucleus sampling with p=0.9p{=}0.9 (Holtzman et al., 2020), and adaptive sampling with $\epsilon{=}0.001$ (Zhu et al., 2024). We evaluate two threshold values $\operatorname{JSD}\leq 0.1$ and $\operatorname{JSD}\leq 0.2$ to study the effect of varying similarity criteria.
@@ 4.3 Results and Discussion
When enforcing stricter similarity requirements ($\operatorname{JSD}\leq 0.1$) we find the distribution becomes flatter and even experiences a U-shaped structure. This suggests that a larger subset of tokens either resolve with short contexts or require nearly the full input to reach distributional convergence. We further analyze this in Appx. D.3 and discover that this bimodality is mainly attributed to the shorter sequences. Indeed, we find that the long documents, for which all sequences have more than 7k tokens, do not experience the bimodality to the same degree.
@@ 5.1 Distributional-aware long-context sequence detection
For concreteness, unless otherwise stated, we fix $\phi$ to nucleus sampling with parameter $p=0.9$ (Holtzman et al., 2020). (Ablation results on different metrics and $p$ values in Appx E.2.) Finally, we fix the prefix length to $32$ based on our findings in Sec. 3 and consistent with Liu et al. (2025).
Controlled validation. To demonstrate that LSDS can effectively detect long-context sequences, we conduct a controlled needle-in-a-haystack experiment adapted from Kamradt (2023). We create short and long queries, such that only in the former the answer appears in the final 32 tokens; see Appx. E.1.1 for details and example illustration. Figure 5 shows the resulting LSDS distributions for Mistral-7B-Instruct (v0.2). We observe a clear separation between categories despite no ground-truth next-token information: short-context sequences concentrate at low LSDS values ($\leq 0.4$), while long-context sequences dominate high LSDS values ($\geq 0.7$). Similar patterns hold for LLaMA-3-8B/Qwen2-7B (see Appx. E.1.1).
@@ 5.2 Evaluation on Natural Text
In natural text, determining the true context dependency requires an oracle with access to the actual next token. We employ two such oracles to establish ground-truth short-context vs long-context labels for sequences. (1) MCL Oracle: For $\mathbf{s}$ with next-token $t$, we classify $\mathbf{s}$ as long-context iff $\mathsf{MCL}\left(\mathbf{s}|t\right)\geq 32$ (see Defn. 1). (2) LSD Oracle: Following Fang et al. (2025), we classify $\mathbf{s}$ as long-context iff $\operatorname{LSD}(\mathbf{s}|t)>2$ & $\operatorname{LCL}(\mathbf{s}|t)\geq-1$ (details in Appx. E.1.4 ).Both oracles require knowledge of ground-truth next token and thus cannot be deployed at inference time. We show that LSDS-based classification agrees well with these oracles despite not knowing the ground-truth.
Figure 6: Distribution of $\mathsf{LSDS}\left(\mathbf{s}\right)$ for short- and long-contexts on GovReport (long doc) and Reddit Writing Prompts (short doc), pooled across three models. Illustrates a clear distinction between long and short contexts.
@@ 6 Long-Context Token Boosting
If so, we could favor generating such tokens when we detect a sequence requires long-context reasoning (which, as shown in the previous section, we can reliably identify). Here, we show this is possible and evaluate our method on text-based question answering, a standard benchmark for assessing long-context reasoning capabilities (Liu et al., 2023; Krishna et al., 2023; Beltagy et al., 2020; Bai et al., 2024).